# Advancing Set-Conditional Set Generation: Graph Diffusion for Fast Simulation of Reconstructed Particles

### v1

@@ 4 Training and Loss
Figure 3: Cardinality distributions for the generated Pflow and target sets for different input set cardinalities.
5 Results
Given the task of set prediction, we can assess multiple aspects of network performance. Firstly, we analyze properties of the set as a whole, such as set cardinality and inclusive particle distributions. Secondly, we examine the properties of individual set constituents. Additionally, for our application, we aim to study resolution modeling, which entails examining the variation in generated particle properties given the same input values. These aspects are discussed separately in the following subsections.
5.1 Set based performance
The cardinality prediction network follows the same principles and structure for both approaches, hence, similar results are expected. Minor deviations may occur as the networks are trained simultaneously with the property prediction network using combined losses. Fig. 3 compares the generated and target set cardinalities for different input set cardinalities. The distributions for low cardinalities peak at the truth cardinality, with the width of the distribution increasing as the number of truth particles rises. Higher particle counts in the event lead to more instances of misreconstructions and inefficiencies. For very high cardinalities, the peak position is lower than the truth cardinality since dense environments are more prone to reconstructing several neutral particles as one. Both networks closely match the target distributions, even for high cardinalities where the event count is low, with predictions aligning with targets within statistical uncertainties. Next, we examine the marginal distributions of the set constituents depicted in Fig. 4. These histograms encompass all particles from all events in the test set. Graph Diffusion outperforms the baseline model in the $p_{T}$ distributions for low energies. Both models perform well in the intermediate and high $p_{T}$ range. The Slot Attention model slightly outperforms Graph Diffusion in the distribution of $\eta$. The bias comes from the fact that low $p_{T}$ particles tend to have higher $|\eta|$ and vice versa. However, the mismatch results in a deviation of a few percent, which is not deemed very significant. The distributions of $\phi$ closely match the target distribution within statistical errors for both networks.
@@ 5.1 Set based performance
The feature prediction task involves classifying particles as charged or neutral. Given the initially predicted total cardinality, the network must learn the number of charged and neutral constituents in the set. Since the total cardinality prediction displays good agreement, we can separately compare the neutral and charged contributions to analyze class prediction performance. Fig. 5 illustrates the difference between the number of truth and reconstructed particles per event, separated for neutral and charged constituents. The target distribution for charged particles is a delta function since track presence ensures perfect classification. However, the reconstruction of neutral particles is more prone to inefficiencies and fakes, resulting in a wider distribution. Slot Attention demonstrates excellent agreement with the target, while Graph Diffusion exhibits some misclassification, as achieving binary prediction accuracy is challenging in diffusion models.
Overall, we conclude that both models reproduce set properties to a very satisfactory extent.
@@
5.2 Constituent based performance
@@
These residual distributions are depicted in Fig. 6. For angular quantities, Graph Diffusion performs better than Slot Attention in matching the peak of the distribution but exhibits wider tails originating from outliers, particularly in $\eta$. Conversely, Slot Attention displays a narrower distribution in $\eta$, suggesting that generated particles are not sufficiently smeared with respect to the truth. This trend is also evident in the right tail of the $p_{T}$ distribution. Here, Graph Diffusion demonstrates better agreement. However, both models deviate from the target distribution in the peak area. It is noteworthy that this area is primarily influenced by low-$p_{T}$ particles, which are challenging to model.
@@
To gain deeper insights, we depict the residual $p_{T}$ distribution for different bins of the matched truth particle $p_{T}$ in Fig. 7. Initially focusing on the target distributions, we observe that low-energy particles form a narrow distribution with sharply declining tails comprised of mismatched or noisy particles. As the truth $p_{T}$ increases, the distributions widen, likely driven by worsening track momentum resolution. Both models tend to overestimate the residuals for low energies. For higher energies, we observe excellent agreement for graph diffusion, while slot attention predicts distributions that are too narrow, underestimating smearing effects and predicting values too close to the truth.
@@
We can summarize previously made observations: Slot Attention exhibits a clear left-shift in the bulk of the data, indicating overly precise predictions. Both networks fail to match the left tail of the distribution, which consists of events with very low energies. In the right peak, we observe a slight shift to the right for the Graph Diffusion. While Graph Diffusion shows room for improvement on the edges of the distributions, it matches the bulk of the data very well and clearly outperforms the baseline model.
@@
5.3 Resolution performance
A crucial aspect for a good detector simulation is accurate resolution modeling. The stochastic nature of particle-material interactions leads to non-trivial smearing of truth particle features. Conducting detector simulation and reconstruction multiple times on the same truth event results in different outcomes each time. To evaluate how well our model reproduces this variability, we introduce a replica dataset, similarly to the concept introduced in [fastsim]. We generate 10,000 unique truth events, and for each truth event, we run detector simulation and reconstruction 100 times. After performing truth matching for all replicas, each truth particle has a matched distribution of particles. For each truth particle, we compute the mean and standard deviation of the matched distribution as a proxy for the detector resolution. This is summarized in bins of truth $p_{T}$ in Fig. 9.
@@ 5.3 Resolution performance
While the means are modeled comparably well by both models, Slot Attention largely underestimates the standard deviations and therefore doesn’t accurately model the resolution. However, Graph Diffusion matches the target resolution significantly better, with a slight trend towards overestimating the resolution.
It’s worth mentioning that utilizing the replica dataset for training with an extended loss function, as demonstrated in [fastsim], is an option. In such a scenario, the Slot Attention model would yield better performance in resolution modelling. However, Graph Diffusion consistently outperforms it. However, since we found no significant advantage for the latter, we opted for the simpler setup without replicas.
For demonstrative purposes, we present an event display with replicas in Fig. 10. The showcased event was selected because it contains particles across a broad $p_{T}$ spectrum without exceeding a count of particles that would classify it as a rare event. The event display aligns with our previous discussions. Slot Attention fails to accurately model the distribution and does not replicate the low-energy noise. In contrast, Graph Diffusion exhibits a much closer alignment with the target distribution but is prone to a few noisy outliers.
6 Conclusion
@@
Central to our approach is the introduction of the Graph Diffusion architecture, which leverages cutting-edge diffusion techniques on graph-valued data. We conducted a comprehensive analysis of the network’s performance, focusing on three key aspects: overall set quantities, features of individual set constituents, and resolution modeling, a significant challenge in detector simulation. Compared to the previous paper, GD allows good resolution modeling without the need for a training dataset containing multiple detector replicas of the same underlying truth event.
To accurately assess resolution performance, we introduced the replica dataset, enabling us to evaluate generation outcomes multiple times per ground truth. Our findings demonstrate that Graph Diffusion significantly outperforms the baseline model in resolution modeling, while exhibiting comparable performance in overall set constituent features.
Looking ahead, our ultimate objective is to develop a fast simulation package that is universally applicable across various events. Future directions could involve scaling up the model to process full events and therefore handle higher cardinalities, although the performance implications of such scaling remain uncertain, particularly for events with several hundred particles. Alternatively, implementing event partitioning could address the cardinality issue, but would require training on a diverse range of detector signatures, including jets of varying energy ranges and isolated objects.
@@ Acknowledgement
ED, EG, NK, DK, and NS are supported by the BSF-NSF grant 2028 and the ISF Research Center 494.
[…]

### v2

@@ IV Training and Loss
Figure 3: Cardinality distributions for the generated PFOs and target sets for different input set cardinalities.
V Results
Given the task of set generation, we can assess multiple aspects of network performance. First, we analyze properties of the set as a whole, such as set cardinality and inclusive particle distributions. Second, we examine the properties of individual set constituents. Additionally, for our application, we aim to study resolution modeling, which entails examining the variation in generated particle properties given the same input values. These aspects are discussed separately in the following subsections.
V.1 Set-based Performance
The cardinality prediction network follows the same principles and structure for all used approaches; hence, similar results are expected. Minor deviations may occur as the networks are trained simultaneously with the property prediction network using combined losses. Figure 3 compares the generated and target set cardinalities for different input set cardinalities. The distributions for low cardinalities peak at the truth cardinality, with the width of the distribution increasing as the number of truth particles rises. Higher particle counts in the event lead to more instances of misreconstructions and inefficiencies. For very high cardinalities, the peak position is lower than the truth cardinality since dense environments are more prone to reconstructing several neutral particles as one. Both networks closely match the target distributions, even for high cardinalities where the event count is low, with predictions aligning with targets within statistical uncertainties. Next, we examine the marginal distributions of the set constituents depicted in Fig. 4. These histograms encompass all particles from all events in the test set. The diffusion models outperform the baseline model in the $p_{T}$ distributions for low energies. Both models perform well in the intermediate and high $p_{T}$ range. The Slot Attention model slightly outperforms the diffusion models in the distribution of $\eta$. The bias comes from the fact that low $p_{T}$ particles tend to have higher $|\eta|$ and vice versa. However, the mismatch results in a deviation of a few percent, which is not deemed very significant. The distributions of $\phi$ closely match the target distribution within statistical errors for both networks.
@@ V.1 Set-based Performance
The feature prediction task involves classifying particles as charged or neutral. Given the initially predicted total cardinality, the network must learn the number of charged and neutral constituents in the set. Since the total cardinality prediction displays good agreement, we can separately compare the neutral and charged contributions to analyze class prediction performance. Figure 5 illustrates the difference between the number of truth and reconstructed particles per event, separated for neutral and charged constituents. The target distribution for charged particles is a delta function since track presence ensures perfect classification. However, the reconstruction of neutral particles is more prone to inefficiencies and fakes, resulting in a wider distribution. Slot Attention demonstrates excellent agreement with the target, while the diffusion models exhibit some misclassification, as achieving binary prediction accuracy is challenging in diffusion models.
Overall, we conclude that both models are able to reproduce the salient properteis of the sets of PFOs.
@@
V.2 Constituent-based Performance
@@
These residual distributions are depicted in Fig. 6. For angular quantities, the diffusion models perform better than Slot Attention in matching the peak of the distribution but exhibits wider tails originating from outliers, particularly in $\eta$. Conversely, Slot Attention displays a narrower distribution in $\eta$, suggesting that generated particles are not sufficiently smeared with respect to the truth. This trend is also evident in the right tail of the $p_{T}$ distribution. Here, the diffusion models demonstrate better agreement. However, EDM deviates from the target distribution in the peak area similarly to the SA, while CFM has a very good match. Notably, this area is primarily influenced by low-$p_{T}$ particles, which are challenging to model.
@@
To gain deeper insights, we depict the residual $p_{T}$ distribution for different bins of the matched truth particle $p_{T}$ in Fig. 7. Initially focusing on the target distributions, we observe that low-energy particles form a narrow distribution with sharply declining tails comprised of mismatched or noisy particles. As the truth $p_{T}$ increases, the distributions widen, likely driven by worsening track momentum resolution. All tend to overestimate the residuals for low energies. For higher energies, we observe excellent agreement for the diffusion models, while slot attention predicts distributions that are too narrow, underestimating smearing effects and predicting values too close to the truth.
@@
We can summarize previously made observations: Slot Attention exhibits a clear left-shift in the bulk of the data, indicating overly precise predictions. Both networks fail to match the left tail of the distribution, which consists of events with very low energies. In the right peak, we observe a slight shift to the right for the diffusion models. While the diffusion models shows room for improvement on the edges of the distributions, they match the bulk of the data very well and clearly outperform the baseline model. The CFM seems somewhat superior to the EDM, in particular for the per-particle Hungarian cost in Fig. 8.
@@
V.3 Resolution Performance
A crucial aspect for a good detector simulation is accurate resolution modeling. The stochastic nature of particle-material interactions leads to non-trivial smearing of truth particle features. Conducting detector simulation and reconstruction multiple times on the same truth event results in different outcomes each time. To evaluate how well our model reproduces this variability, we introduce a replica dataset, similarly to the concept introduced in Ref. [38]. We generate 10,000 unique truth events, and for each truth event, we run detector simulation and reconstruction 100 times. After performing truth matching for all replicas, each truth particle has a matched distribution of particles. For each truth particle, we compute the mean and standard deviation of the matched distribution as a proxy for the detector resolution. This is summarized in bins of truth $p_{T}$ in Fig. 9.
@@ V.3 Resolution Performance
While the means are modeled comparably well by both models, Slot Attention largely underestimates the standard deviations and therefore doesn’t accurately model the resolution. However, the diffusion models match the target resolution significantly better. EDM features a slight trend towards underestimating the resolution.
We note that utilizing the replica dataset for training with an extended loss function, as demonstrated in Ref. [38], is an option. In such a scenario, the Slot Attention model would yield better performance in resolution modelling. However, the diffusion models consistently outperform it. Since we found no significant advantage for the latter, we opted for the simpler setup without replicas.
For demonstrative purposes, we present an event display with replicas in Fig. 10. The showcased event was selected because it contains particles across a broad $p_{T}$ spectrum without exceeding a count of particles that would classify it as a rare event. The event display aligns with our previous discussions. Slot Attention fails to accurately model the distribution and does not replicate the low-energy noise. In contrast, the diffusion models exhibit a much closer alignment with the target distribution but are prone to a few noisy outliers.
VI Conclusion
@@
Central to our approach is the introduction of the diffusion models, which leverage cutting-edge diffusion techniques on graph-valued data. We conducted a comprehensive analysis of the networks’ performance, focusing on three key aspects: overall set quantities, features of individual set constituents, and resolution modeling, a significant challenge in detector simulation. Both diffusion models outperform slot attention with respect to resolution modelling without the need for a training dataset containing multiple detector replicas of the same underlying truth event. Overall, the CFM model outperformed the EDM, seen most clearly for the per-event Hungarian matching cost (Fig. 8).
To accurately assess resolution performance, we introduced the replica dataset, enabling us to evaluate generation outcomes multiple times per ground truth. Our findings demonstrate that the diffusion models significantly outperform the baseline model in resolution modelling while exhibiting comparable performance in overall set constituent features.
Looking ahead, our ultimate objective is to develop a fast simulation package that is universally applicable across various events. Future directions could involve scaling up the model to process full events and, therefore, handle higher cardinalities, although the performance implications of such scaling remain uncertain, particularly for events with several hundred particles. Alternatively, implementing event partitioning could address the cardinality issue, but would require training on a diverse range of detector signatures, including jets of varying energy ranges and isolated objects. Using the results of this paper, Ref. [40] makes progress on some of these challenges towards a full end-to-end approach.
Data and Code Availability
The code for this paper can be found at https://github.com/dkobylianskii/fastsim-advanced. The COCOA dataset can be found at Zenodo at https://zenodo.org/records/11383444.
@@ Acknowledgement
We thank Yaron Lipman and Vinicius Mikuni for fruitful discussions. ED, EG, NK, DK, and NS are supported by the BSF-NSF grant 2028 and the ISF Research Center 494. BN is supported by the U.S. Department of Energy (DOE), Office of Science under contract DE-AC02-05CH11231.